in an economy investment exceeds private saving by RS 10,000, budget d if any wrong with the data?eficit is RS 6,000 and net exports is RS 4,000. What if any wrong with the data?
1
Expert's answer
2020-04-06T10:08:40-0400

National saving is "S = Y - C - G = NX + I = Sg + Sp," where Sg is public saving, Sp is private saving, C is consumption, I is investment, Y is GDP, G is government purchases, NX is net exports.

Sg = -6,000,

I - Sp = 10,000,

NX = 4,000, so:

4,000 + 10,000 = - 6,000, which is incorrect.